Give a rough estimate (by rounding off to nearest hundreds) and also a closer estimate (by rounding off to nearest tens): 
439 + 334 + 4317 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

439 + 334 + 4317
Rounding off to nearest hundreds
439 + 334 + 4317 = 400 + 300 + 4300 = 5000
Rounding off to nearest tens
439 + 334 + 4317 = 440 + 330 + 4320 = 5090
​​
Thanks